# Cubic power functions with optimal second-order differential uniformity

### v1

@@ Abstract.
We discuss the second-order differential uniformity of vectorial Boolean functions, a relevant cryptographic property due to indication of resistance to the boomerang attack. First, we discuss connections with the second-order zero differential uniformity and its recent literature. We then prove the optimality of monomial functions with univariate form $x^{d}$ where $d=2^{2k}+2^{k}+1$ and $\gcd(k,n)=1$, and begin work towards generalising such conditions to all monomial functions of algebraic degree 3. Finally, we discuss further questions arising from computational results.
@@ 1. Introduction and background
Boolean functions can be understood as mathematical functions that model transformations on bit-strings, and have wide usage across theoretical computer science, primarily in the subjects of coding theory, and cryptography. Extensive coverage of the application of Boolean functions in cryptography and coding theory is provided in [carlet2021book, cusick2009cryptographic]. Boolean functions have historically been used in the construction of stream ciphers, pseudorandom generators, and block ciphers. In particular, the S-boxes in the Data Encryption Standard (DES) and Advanced Encryption Standard (AES) ciphers can be described using Boolean functions [carlet2021book]. The difficulty lies in choosing a “secure” Boolean function (or multiple) in the construction of these ciphers in order to provide resistance against attacks.
One well-studied attack on cryptosystems is the differential attack, introduced by [biham1991differential]. For a Boolean function to be resistant to a differential attack, it must satisfy a criterion equivalent to the outputs of the derivative $D_{a}f$ being as uniformly distributed as possible [carlet2021book]*Ch. 3.4.1, or equivalently, having minimal differential uniformity [nyberg1993differential, nyberg1993provable]. This motivates the extensive study of perfect nonlinear and APN functions, as explained in Section 2.
@@
The main direction to-date in studying resistance to this attack is to generalise the differential uniformity to the second-order derivative of the chosen Boolean function. In [cid2018bct], the Boomerang Connectivity Table (BCT) was introduced to measure the resistance of an S-box to the differential attack. However, this table does not address the case of Feistel ciphers, and thus later, in [boukerrou2020feistelboomerang], the Feistel Boomerang Connectivity Table (FBCT) was introduced for this purpose. More recently, there has been a focus on the second-order zero differential spectra and the second-order zero differential uniformity, which is defined as follows. We note that the authors of [li2022secondorder] originally call this the second-order differential spectra, but we include the term ”zero” to avoid confusion with the second-order differential uniformity we work with in later sections.
Definition 1.1 ([li2022secondorder]).
$\nabla_{f}=\max\{\nabla_{f}(a,b)\ :\ a\neq b,\ a,b\neq 0\}.$
$\nabla_{f}=\max\{\nabla_{f}(a,b)\ :\ a,b\neq 0\}.$
@@
This notion extends the Feistel Boomerang Connectivity Table to Boolean functions over odd characteristic fields; indeed the definitions of the Feistel Boomerang uniformity and the second-order zero differential uniformity are identical for $p=2$. There have been a number of results on the second-order zero differential spectra of various classes of functions. We provide a full list of studied power functions in Table 1, along with their second-order zero differential uniformity, denoted by $\nabla_{f}$. We note that, while we do not list them, full descriptions of the second-order zero differential spectra for each table entry can be found in the respective citations provided.
$p$ $d$ Condition $\nabla_{f}$ Reference
$p=2$ $22^{n}-$ 2 $n$ odd, or $n$ even 0, or 4 [eddahmani2022explicit]*Thm. 10
$p=2$ $12^{k}+$ 1 $\gcd(k,n)=m$ 00 [eddahmani2022explicit]*Thm. 12
$p=2$ $12^{2k}+2^{k}+$ 1 $n=2^{4k}$ $2^{2k}$ [eddahmani2022explicit]*Thm. 14
$p=2$ $12^{m+1}-$ 1 $n=2m+1$ or $n=2m$ 0 or $2^{m}$ [man2023indepth]*Cor. 1
$p=2$ $12^{m}-$ 1 $n=2m+1$ or $n=2m$ $42^{m}-$ 4 [garg2023secondorder]*Thm. 4.1
$p=2$ 21 $n$ odd 4 [garg2023secondorderapn]*Thm 4.1
$p=2$ $2^{n}-2^{s}$ $\gcd(n,s+1)=1$, $n-s=3$ 4 [garg2023secondorderapn]*Thm. 4.9
$p=2$ 7 any 4 [man2023secondorder]*Thm. 1
$p=2$ $32^{m+1}+$ 3 $n=2m+1$, or $n=2m$ 4 or $2^{m}$ [man2023secondorder]*Thm. 2
$p=3$ $33^{n}-$ 3 $n>1$ odd 2 [li2022secondorder]*Thm. 3.2
$p=3$ $23^{n}-$ 2 any 3 [li2022secondorder]*Thm. 3.3
$p=3$ $\frac{3^{n}-1}{2}+2$ $n$ odd 3 [garg2023secondorder]*Thm. 3.5
$p=3$ $12\cdot\frac{3^{n}-1}{2}+$ 1 any 3 [garg2023secondorderapn]*Thm 3.8
$p=3$ 5 any 3 [man2023secondorder]*Thm. 3
$p=3$ 7 any 3 [man2023secondorder]*Thm. 4
$p>3$ 3 any 1 [li2022secondorder]*Thm. 3.1
[…]
$p>3$ 4 $n>1$ 2 [garg2023secondorder]*Prop. 3.4
$p>3$ $\frac{p^{k}+1}{2}$ $\gcd(2n,k)=1$ $\frac{p-3}{2}$ [garg2023secondorder]*Th. 3.2
[…]
$p>5$ 5 any 3 [man2023secondorder]*Thm. 3
@@
Table 1. Power functions $f(x)=x^{d}$ over $\mathbb{F}_{p^{n}}$ with known second-order zero differential uniformity, a corrected and updated form of [man2023secondorder]*Table 1.
Our approach follows the definition of the second-order differential uniformity provided in [aubry2018differential], which we feel most naturally extends the definition of the differential uniformity. Our key result provides a new class of power functions with optimal second-order differential uniformity, namely the functions $f:\mathbb{F}_{2^{n}}\to\mathbb{F}_{2^{n}}$ given by $f(x)=x^{d}$, where $d=2^{2k}+2^{k}+1$, and $\gcd(k,n)=1$. We note that the second-order zero differential uniformity and second-order differential uniformity are equivalent for cubic Boolean functions, as the second derivatives appearing in both cases are affine. As such, many of the studied functions listed in Table 1 are relevant here, and our results are directly relevant to the study of the second-order zero differential uniformity.
@@
This paper is structured as follows. In Section 2, we discuss preliminaries of polynomials in finite fields, Boolean functions, and differential uniformity. In Section 3, we discuss a class of power functions of algebraic degree 3 and provide explicit values of the second-order differential uniformity in all cases. In Section 4, we discuss algebraic degree 3 power functions in the general case, and provide conditions for optimal second-order differential uniformity. Finally, we conclude in Section 5 by discussing further questions arising from computational results.
@@ 2. Preliminaries
We use $p$ to denote a prime, and $q$ to denote a prime power. We then denote by $\mathbb{F}_{q}$ the finite field containing $q$ elements, and by $\mathbb{F}_{q}$ the non-zero elements of $\mathbb{F}_{q}^{*}$. Most commonly, we choose $p=2$, $q=2^{n}$ for some $n\in\mathbb{N}$, and write $\mathbb{F}_{2^{n}}$. Similarly, we denote by $\mathbb{F}_{p}^{n}$ the vector space of dimension $n$ over the field $\mathbb{F}_{p}$ and, most commonly choosing $p=2$, write $\mathbb{F}_{2}^{n}$. We use the convention =$\mathbb{N}=\{1,2,3,\dots\}$, i.e. not including 00.
@@
Given two Boolean functions $f,g:\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{m}$, we say that $f,g$ are affine equivalent if there exists some affine automorphims $L_{1},L_{2}$ of $\mathbb{F}_{2}^{n},\mathbb{F}_{2}^{m}$ respectively, such that $f=L_{2}\circ g\circ L_{1}$.
@@
$f(x)=\sum_{i=0}^{2^{n}-1}\delta_{i}x^{i}$ (2.2)
where $\delta_{i}\in\mathbb{F}_{2^{n}}$, via the isomorphism between $\mathbb{F}_{2}^{n}$ and $\mathbb{F}_{2^{n}}$.
When we write a Boolean function in the form of 2.1, we say it is in multivariate form, and when we write it in the form of 2.2, we say it is in univariate form. We will primarily work with Boolean functions in univariate form from here.
Linearised polynomials
Due to [garg2023secondorderapn]*Thm. 4.3, linearised polynomials form a key element in the study of the second-order differential uniformity of Boolean functions.
Definition 2.5 ([lidl1994introduction]*Def. 3.49).
A polynomial of the form:
$L(x)=\sum_{i=0}^{m}\alpha_{i}x^{p^{i}},$
where $\alpha_{i}\in\mathbb{F}_{p^{n}}$ and $p$ a prime, is called a $p$-polynomial over $\mathbb{F}_{p^{n}}$. If the value of $p$ is fixed or clear from context, $L$ may also be called a linearised polynomial.
Note that the term linearised comes from the fact that the polynomial $L$ can be considered as a linear operator over $\mathbb{F}_{p^{n}}$.
Theorem 2.6 ([lidl1994introduction]*Thm. 1.46).
Let $R$ be a commutative ring of characteristic $p$. Then for all $a,b\in R$:
$(a+b)^{p^{n}}=a^{p^{n}}+b^{p^{n}},\textrm{and }(a-b)^{p^{n}}=a^{p^{n}}-b^{p^{n}}$
@@
We briefly mention the paper [aubry2018differential], which studies the distribution of the second-order derivative of polynomials over $\mathbb{F}_{2^{n}}$. They first define the second-order differential uniformity as:
$\delta^{2}(f)=\max_{a\neq a^{\prime}\in\mathbb{F}_{2^{n}}^{*},\ b\in F_{2^{n}}}\lvert\{x\in\mathbb{F}_{2^{n}}\ :\ D^{(2)}_{a,a^{\prime}}f(x)=b\}\rvert.$
Note that an equivalent definition is given recently in [eddahmani2024doublediff], referred to as the double differential uniformity. The definition we use in the following work matches these definitions, with altered notation.
@@
Given $n\in\mathbb{N}$, let $f:\mathbb{F}_{2^{n}}\to\mathbb{F}_{2^{n}}$ be a Boolean function. Then we say that $f$ has second-order differential uniformity $\nabla_{f}$ if $\nabla_{f}$ is the least integer such that, for every $a,b\in\mathbb{F}_{2^{n}}^{*}$, $a\neq b$, and every $c\in\mathbb{F}_{2^{n}}$, the equation $D_{a,b}^{(2)}f(x)=c$ has at most $\nabla_{f}$ solutions, i.e.,
$\nabla_{f}=\max_{a\neq b\in\mathbb{F}_{2^{n}}^{*},\ c\in F_{2^{n}}}\lvert\{x\in\mathbb{F}_{2^{n}}\ :\ D^{(2)}_{a,b}f(x)=c\}\rvert.$
@@
and thus for any Boolean function $f$, we have $\nabla_{f}\geq 4$. Moreover, $\nabla_{f}$ is always a multiple of 4.
@@
Given $n\in\mathbb{N}$, let $f:\mathbb{F}_{2^{n}}\to\mathbb{F}_{2^{n}}$ be a Boolean function. We say that $f$ has optimal second-order differential uniformity if $\nabla_{f}=4$.
Note that the second-order differential uniformity is an affine invariant.
@@ 3. Power functions with exponent $12^{2k}+2^{k}+$ 1
By [eddahmani2022explicit]*Prop. 7, it is equivalent to consider only the derivatives $D_{1,c}^{(2)}f$, where $c\in\mathbb{F}_{2^{n}}^{*}$. Let $q=2^{k}$. We then have:
@@
Clearly, we have solutions for $z\in\{0,1\}$, so assume $z\notin\{0,1\}$. As $c\notin\{0,1\}$ and $\gcd(k,n)=1$, we have $c^{q^{2}}\neq c^{q}$, $c^{q^{2}}\neq c$, and $c^{q}\neq c$. We then have:
@@
Dividing 3.1 by $(c^{q}+c)$, from 3.2 and 3.3 we have:
@@
which implies $z+c\in\{0,1\}$. Thus the only solutions are $z\in\{0,1,c,1+c\}$, so $D_{1,c}^{(2)}f$ is 4-to-1. As this holds for all $c\in\mathbb{F}_{2^{n}}^{*}$, we have $\nabla_{f}=4$. ∎
@@
From 3.1, we have:
00 0 $=z^{q^{2}}(c^{q}+c)+z^{q}(c^{q^{2}}+c)+z(c^{q^{2}}+c^{q}).$
where $q=2^{k}$. As $\gcd(k,n)=u>1$, we have that $\mathbb{F}_{2^{u}}$ is a subfield of $\mathbb{F}_{2^{n}}$. Thus whenever $c\in\mathbb{F}_{2^{u}}\setminus\{0,1\}$ we have $c^{q^{2}}=c^{q}=c$, so choosing any such $c$ we have that 3.1 holds for all $z$. As such, $D_{1,c}^{(2)}f(0)=D_{1,c}^{(2)}f(0+z)$ for all $z\in\mathbb{F}_{2^{n}}$, so $f$ has second-order differential uniformity $2^{n}$. ∎
Combining the above, we find the following result.
@@
Proof.
By Proposition 3.2, we have that $\gcd(k,n)>1$ implies that $f$ is not optimal, and so by the contrapositive, we have that $f$ being optimal implies $\gcd(k,n)=1$. Along with Theorem 3.1, this completes the if and only if statement. The second statement is exactly Proposition 3.2. ∎
Note that $f$ having second-order uniformity $2^{n}$ does not provide information on the exact value of $\gcd(k,n)$.

### v2

@@ Abstract.
We discuss the second-order differential uniformity of vectorial Boolean functions. The closely related notion of second-order zero differential uniformity has recently been studied in connection to resistance to the boomerang attack. We prove that monomial functions with univariate form $x^{d}$ where $d=2^{2k}+2^{k}+1$ and $\gcd(k,n)=1$ have optimal second-order differential uniformity. Computational results suggest that, up to affine equivalence, these might be the only optimal cubic power functions. We begin work towards generalising such conditions to all monomial functions of algebraic degree 3. We also discuss further questions arising from computational results.
@@ 1. Introduction and background
Boolean functions can be understood as mathematical functions that model transformations on bit-strings, and have wide usage across theoretical computer science; we are primarily considering cryptography and coding theory. Extensive coverage of the application of Boolean functions in cryptography and coding theory is provided in [carlet2021book, cusick2009cryptographic]. Boolean functions have historically been used in symmetric cryptography, for the construction of both stream ciphers and block ciphers. In particular, the S-boxes in the Data Encryption Standard (DES) and Advanced Encryption Standard (AES) ciphers can be described using Boolean functions. The difficulty lies in choosing “secure” Boolean functions in the construction of these ciphers in order to provide resistance against attacks.
One well-studied attack on cryptosystems is the differential attack, introduced by [biham1991differential]. For a Boolean function to be resistant to a differential attack, it must satisfy a criterion equivalent to the outputs of the derivative $D_{a}f$ being as uniformly distributed as possible [carlet2021book]*Ch. 3.4.1, or equivalently, having minimal differential uniformity [nyberg1993differential, nyberg1993provable]. This motivates the extensive study of functions with optimal differential uniformity, namely perfect nonlinear and APN functions, as explained in Section 2.
@@
The main direction to-date in studying resistance to this attack is to generalise the differential uniformity to the second-order derivative of the chosen Boolean function. In [cid2018bct], the Boomerang Connectivity Table (BCT) was introduced to measure the resistance of an S-box to the differential attack. However, this table does not address the case of Feistel ciphers, and thus later, in [boukerrou2020feistelboomerang], the Feistel Boomerang Connectivity Table (FBCT) was introduced for this purpose. More recently, there has been a focus on the second-order zero differential spectra and the second-order zero differential uniformity. This notion extends the Feistel Boomerang Connectivity Table to Boolean functions over odd characteristic fields; indeed the definitions of the Feistel Boomerang uniformity and the second-order zero differential uniformity are identical for $p=2$. There have been a number of results on the second-order zero differential spectra of various classes of functions. We provide a full list of studied power functions over $\mathbb{F}_{2^{n}}$ in Table 1, along with their second-order zero differential uniformity, denoted by $\nabla_{f}$. We note that, while we do not list them, full descriptions of the second-order zero differential spectra for each table entry can be found in the respective citations provided.
$d$ Condition $\nabla_{f}$ Reference
$22^{n}-$ 2 $n$ odd, or $n$ even 0, or 4 [eddahmani2022explicit]*Thm. 10
$12^{k}+$ 1 $\gcd(k,n)=1$ 00 [eddahmani2022explicit]*Thm. 12
$12^{2k}+2^{k}+$ 1 $n=2^{4k}$ $2^{2k}$ [eddahmani2022explicit]*Thm. 14
$12^{m+1}-$ 1 $n=2m+1$ or $n=2m$ 0 or $2^{m}$ [man2023indepth]*Cor. 1
$12^{m}-$ 1 $n=2m+1$ or $n=2m$ $42^{m}-$ 4 [garg2023secondorder]*Thm. 4.1
21 $n$ odd 4 [garg2023secondorderapn]*Thm 4.1
$2^{n}-2^{s}$ $\gcd(n,s+1)=1$, $n-s=3$ 4 [garg2023secondorderapn]*Thm. 4.9
7 any 4 [man2023secondorder]*Thm. 1
$32^{m+1}+$ 3 $n=2m+1$, or $n=2m$ 4 or $2^{m}$ [man2023secondorder]*Thm. 2
$12^{2k}+2^{k}+$ 1 $\gcd(k,n)=1$ $\leq 4$ This paper, Cor. 3.4
@@
Table 1. Power functions $f(x)=x^{d}$ over $\mathbb{F}_{2^{n}}$ with known second-order zero differential uniformity.
In [aubry2018differential], the authors introduce the second-order differential uniformity, denoted by $\delta^{2}(f)$, and compute the second-order differential uniformity of the inverse function, see Table 2. Our approach follows this definition, which we feel most naturally extends the definition of the differential uniformity. Note that $\delta^{2}(f)\geq\nabla_{f}$, so any results on the second-order differential uniformity also yield an upper bound on the second-order zero differential uniformity. We also mention the paper [eddahmani2024doublediff], where the authors define the double differential uniformity, a similar notion defined for Boolean functions from $\mathbb{F}_{2^{n}}$ to $\mathbb{F}_{2^{m}}$. The authors then compute values of the double differential uniformity for the inverse, Gold, and Bracken-Leander functions. When $n=m$, the definition of the double differential uniformity is identical to that of the second-order differential uniformity, and thus these results are directly relevant to our work. We summarise results on functions of known second-order differential uniformity, denoted by $\delta^{2}(f)$, in Table 2.
$d$ Condition $\delta^{2}(f)$ Reference
$22^{n}-$ 2 $n\geq 6$ 8 [aubry2018differential]*Prop. 8.1
$12^{k}+$ 1 $\gcd(k,n)=1$ $2^{n}$ [eddahmani2024doublediff]*Cor. 2
$12^{2k}+2^{k}+$ 1 $n=2^{4k}$ $2^{n}$ [eddahmani2024doublediff]*Cor. 3
$12^{2k}+2^{k}+$ 1 $\gcd(k,n)=1$ $4$ This paper, Thm. 3.1
$12^{2k}+2^{k}+$ 1 $\gcd(k,n)>1$ $2^{n}$ This paper, Prop. 3.2
Table 2. Power functions $f(x)=x^{d}$ over $\mathbb{F}_{p^{n}}$ with known second-order differential uniformity
@@
Our key result, Theorem 3.1, provides a new class of power functions with optimal second-order differential uniformity (the optimal value being 4), namely the functions $f:\mathbb{F}_{2^{n}}\to\mathbb{F}_{2^{n}}$ given by $f(x)=x^{d}$, where $d=2^{2k}+2^{k}+1$, and $\gcd(k,n)=1$. This also gives an upper bound on their second-order zero differential uniformity We also discuss algebraic degree 3 power functions in the general case, and provide necessary conditions for optimal second-order differential uniformity (see Section 4). Finally, we conclude in Section 5 by discussing further questions arising from computational results and conjecture that the cubic functions described in Theorem 3.1 are, up to affine equivalence, the only cubic power functions with optimal second-order differential uniformity.
@@ 2. Preliminaries
We use $p$ to denote a prime, and $q$ to denote a prime power. We then denote by $\mathbb{F}_{q}$ the finite field containing $q$ elements, and by $\mathbb{F}_{q}^{*}$ the non-zero elements of $\mathbb{F}_{q}$. Most commonly, we choose $p=2$, $q=2^{n}$ for some $n\in\mathbb{N}$, and write $\mathbb{F}_{2^{n}}$. Similarly, we denote by $\mathbb{F}_{p}^{n}$ the vector space of dimension $n$ over the field $\mathbb{F}_{p}$ and, most commonly choosing $p=2$, write $\mathbb{F}_{2}^{n}$. We use the convention =$\mathbb{N}=\{1,2,3,\dots\}$, i.e. not including 00.
@@
Given Boolean functions $f,g:\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}^{m}$, we say that $f,g$ are affine equivalent if there exists some affine automorphims $L_{1},L_{2}$ of $\mathbb{F}_{2}^{n},\mathbb{F}_{2}^{m}$ respectively, such that $f=L_{2}\circ g\circ L_{1}$.
@@
$f(x)=\sum_{i=0}^{2^{n}-1}b_{i}x^{i}$ (2.2)
where $b_{i}\in\mathbb{F}_{2^{n}}$, via the isomorphism between $\mathbb{F}_{2}^{n}$ and $\mathbb{F}_{2^{n}}$.
When we write a Boolean function in the form of (2.1), we say it is in multivariate form, and when we write it in the form of (2.2), we say it is in univariate form. We will primarily work with Boolean functions in univariate form from here.
@@
We then define the second-order differential uniformity, following from [aubry2018differential] with slight notation changes.
@@
Given $n\in\mathbb{N}$, let $f:\mathbb{F}_{2^{n}}\to\mathbb{F}_{2^{n}}$ be a Boolean function. Then we say that $f$ has second-order differential uniformity $\delta^{2}(f)$ if $\delta^{2}(f)$ is the least integer such that, for every $a,b\in\mathbb{F}_{2^{n}}^{*}$, $a\neq b$, and every $c\in\mathbb{F}_{2^{n}}$, the equation $D_{a,b}^{(2)}f(x)=c$ has at most $\delta^{2}(f)$ solutions, i.e.,
$\delta^{2}(f)=\max_{a\neq b\in\mathbb{F}_{2^{n}}^{*},\ c\in F_{2^{n}}}\lvert\{x\in\mathbb{F}_{2^{n}}\ :\ D^{(2)}_{a,b}f(x)=c\}\rvert.$
@@
and thus for any Boolean function $f$, we have $\delta^{2}(f)\geq 4$. Moreover, this gives that $\delta^{2}(f)$ is always a multiple of 4.
@@
Given $n\in\mathbb{N}$, let $f:\mathbb{F}_{2^{n}}\to\mathbb{F}_{2^{n}}$ be a Boolean function. We say that $f$ has optimal second-order differential uniformity if $\delta^{2}(f)=4$.
Note that the second-order differential uniformity is an (extended) affine invariant [eddahmani2024doublediff]*Prop. 2.
The following observation is used widely, e.g. in the proof of [aubry2018differential]*Prop. 8.1. It can also be considered as a corollary of [eddahmani2024doublediff]*Prop. 3. We state it here for later reference.
Proposition 2.11.
Let $f:\mathbb{F}_{2^{n}}\to\mathbb{F}_{2^{n}}$ be a power function defined by $f(x)=x^{d}$. Then, for any $a,b\in\mathbb{F}_{2^{n}}^{*}$ we have $D_{a,b}f(x)=a^{d}D_{1,\frac{b}{a}}f\left(\frac{x}{a}\right)$. Therefore the second-order differential uniformity of $f$ is given by:
$\delta^{2}(f)=\max_{c\in\mathbb{F}_{2^{n}}\setminus\{0,1\},\ c^{\prime}\in F_{2^{n}}}\lvert\{x\in\mathbb{F}_{2^{n}}\ :\ D^{(2)}_{1,c}f(x)=c^{\prime}\}\rvert.$
Proof.
Given $a,b\in\mathbb{F}_{2^{n}}^{*}$, with $a\neq b$, we compute the second derivative $D_{a,b}f$.
$D_{a,b}f(x)$ $=x^{d}+(x+a)^{d}+(x+b)^{d}+(x+a+b)^{d}$
$=a^{d}\left(\left(\frac{x}{a}\right)^{d}+\left(\frac{x}{a}+1\right)^{d}+\left(\frac{x}{a}+\frac{b}{a}\right)^{d}+\left(\frac{x}{a}+1+\frac{b}{a}\right)^{d}\right)$
$=a^{d}D_{1,\frac{b}{a}}f\left(\frac{x}{a}\right).$
Relabeling $y=\frac{x}{a}$, and $c=\frac{b}{a}$, we get that for any $c^{\prime}\in\mathbb{F}_{2^{n}}$:
|$\lvert\{x\in\mathbb{F}_{2^{n}}\ :\ D^{(2)}_{a,b}f(x)=c^{\prime}\}\rvert=\lvert\{y\in\mathbb{F}_{2^{n}}\ :\ D^{(2)}_{1,c}f(y)=a^{-d}c^{\prime}\}\rvert.$
Finally, we define the second-order zero differential uniformity.
Definition 2.12 ([li2022secondorder]).
$\nabla_{f}=\max\{\nabla_{f}(a,b)\ :\ a\neq b,\ a,b\neq 0\}.$
$\nabla_{f}=\max\{\nabla_{f}(a,b)\ :\ a,b\neq 0\}.$
Note that when $p=2$ we always have $\nabla_{f}\leq\delta^{2}(f)$, and similarly to $\delta^{2}(f)$, $\nabla_{f}$ is always a multiple of 4.
@@ 3. Power functions with exponent $12^{2k}+2^{k}+$ 1
By Proposition 2.11, it is equivalent to consider only the derivatives $D_{1,c}^{(2)}f$, where $c\in\mathbb{F}_{2^{n}}\setminus\{0,1\}$. Let $q=2^{k}$. We then have:
@@
Clearly, we have solutions of (3.2) for $z\in\{0,1\}$, so assume $z\notin\{0,1\}$. As $c\notin\{0,1\}$ and $\gcd(k,n)=1$, we have $c^{q^{2}}\neq c^{q}$, $c^{q^{2}}\neq c$, and $c^{q}\neq c$. We then have:
@@
Dividing (3.2) by $(c^{q}+c)$, from (3.3) and (3.4) we have:
@@
which implies $z+c\in\mathbb{F}_{2^{k}}\cap\mathbb{F}_{2^{n}}$. Since $\gcd(k,n)=1$, we have $\mathbb{F}_{2^{k}}\cap\mathbb{F}_{2^{n}}=\mathbb{F}_{2}$. Thus the only solutions of (3.2) are $z\in\{0,1,c,1+c\}$, so $D_{1,c}^{(2)}f$ is 4-to-1. As this holds for all $c\in\mathbb{F}_{2^{n}}^{*}$, we have $\delta^{2}(f)=4$. ∎
@@
Using the notation from the proof of Theorem 3.1, we have from (3.2) that:
00 0 $=z^{q^{2}}(c^{q}+c)+z^{q}(c^{q^{2}}+c)+z(c^{q^{2}}+c^{q}),$
where $q=2^{k}$. As $\gcd(k,n)=u>1$, we have that $\mathbb{F}_{2^{u}}$ is a subfield of $\mathbb{F}_{2^{n}}$. Thus whenever $c\in\mathbb{F}_{2^{u}}\setminus\{0,1\}$ we have $c^{q^{2}}=c^{q}=c$, so choosing any such $c$ we have that (3.2) holds for all $z$. As such, $D_{1,c}^{(2)}f(0)=D_{1,c}^{(2)}f(0+z)$ for all $z\in\mathbb{F}_{2^{n}}$, so $f$ has second-order differential uniformity $2^{n}$. ∎
Note that $f$ having second-order uniformity $2^{n}$ does not provide information on the exact value of $\gcd(k,n)$. Discussing Table 2, Proposition 3.2 contains and expands on the known second-order differential uniformity of the Bracken-Leander function (row 3).
Combining the two results above, we find the following result.
@@
Note that as $\nabla_{f}\leq\delta^{2}(f)$, we also have the following result.
Corollary 3.4.
Let $k,n\in\mathbb{N}$ satisfy $\gcd(k,n)=1$. Then the Boolean function $f:\mathbb{F}_{2^{n}}\to\mathbb{F}_{2^{n}}$ given by $f(x)=x^{2^{2k}+2^{k}+1}$ has second-order zero differential uniformity $\nabla_{f}\leq 4$.